# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ 3.1 Defining the form of the posterior
To learn the posterior distribution $P^{{\bm{z}}|{\bm{x}}}$ in context, we use the framework of flow matching (Lipman et al., 2022). More specifically, we utilize continuous normalizing flows (CNFs) to specify and ultimately sample from $P^{{\bm{z}}|{\bm{x}}}$. CNFs, currently excelling in the field of image synthesis (Esser et al., 2024), do not only allow to flexibly learn almost arbitrary distributions, but are also found to be more sample-efficient in training than for instance diffusion objectives (Lipman et al., 2022; Wildberger et al., 2024). Furthermore, unlike discrete normalizing flows (Papamakarios et al., 2021a), CNF objectives do not limit the architecture of the used neural network, allowing to incorporate complex conditioning on the data ${\bm{x}}$ in addition to flexibly modeling the posterior, which is a crucial aspect of our ICL framework. Refer to Appendix K for more information on CNFs.
@@ 3.1.1 Normalizing flows
Therefore, one assumes that samples from $P^{{\bm{z}}|{\bm{x}}}$ are generated by first drawing ${\bm{z}}_{0}\sim P_{\mathcal{B}}$, and then applying $\psi_{\theta}(\cdot|{\bm{x}})$, such that $\psi_{\theta}({\bm{z}}_{0}|{\bm{x}})\sim P^{{\bm{z}}|{\bm{x}}}$. The base distribution $P_{\mathcal{B}}$ is commonly set to be a standard normal distribution, i.e., $P_{\mathcal{B}}=\mathcal{N}(0,I)$. The conditional flow $\psi_{\theta}(\cdot|{\bm{x}})$ is the object to be learned, such that our model of $P^{{\bm{z}}|{\bm{x}}}$ is defined as $Q^{{\bm{z}}|{\bm{x}}}_{\theta}\vcentcolon=[\psi_{\theta}(\cdot|{\bm{x}})]_{\sharp}P_{\mathcal{B}}$.
@@ 3.1.2 Continuous normalizing flows
$\frac{d}{dt}\psi_{\theta,t}({\bm{z}}|{\bm{x}})=v^{\theta}_{t,{\bm{x}}}(\psi_{\theta,t}({\bm{z}}|{\bm{x}})),\hskip 28.45274pt\psi_{\theta,0}({\bm{z}}|{\bm{x}})={\bm{z}},$ (4)
where $0\leq t\leq 1$. The first condition $\frac{d}{dt}\psi_{\theta,t}({\bm{z}}|{\bm{x}})=v^{\theta}_{t,{\bm{x}}}(\psi_{\theta,t}({\bm{z}}|{\bm{x}}))$ means that $v^{\theta}_{t,{\bm{x}}}$ describes the change in $\psi_{\theta,t}({\bm{z}}|{\bm{x}})$ at time $t$, and the second condition $\psi_{\theta,0}({\bm{z}}|{\bm{x}})={\bm{z}}$ implies that initially the flow is just the identity. The family of vector fields $v^{\theta}_{t,{\bm{x}}}$ is parameterized by a neural network whose parameters $\theta$ will be learned. In order to ultimately compute the flow $v^{\theta}_{1,{\bm{x}}}$, that yields $Q^{{\bm{z}}|{\bm{x}}}_{\theta}=[\psi_{\theta,1}(\cdot|{\bm{x}})]_{\sharp}P_{\mathcal{B}}$, a numerical ODE solver can be used to forward-solve the ODE, which ultimately corresponds to evaluating $\psi_{1,{\bm{x}}}$ at a datapoint ${\bm{z}}_{0}\sim P_{\mathcal{B}}$.
@@
$d_{\textrm{CFM}}\left(Q^{{\bm{z}}|{\bm{x}}}_{\theta},P^{{\bm{z}}|{\bm{x}}}\right)\vcentcolon=\mathbb{E}\left[\left\lvert\left\rvert v_{t,{\bm{x}}}^{\theta}((1-(1-\sigma_{\textrm{min}})t){\bm{z}}_{0}+t{\bm{z}}_{1})-\frac{{\bm{z}}_{1}-(1-\sigma_{\textrm{min}}){\bm{z}}_{0}}{1-(1-\sigma_{\textrm{min}})t}\right\lvert\right\rvert_{2}^{2}\right],$ (5)
where the expectation is taken w.r.t. to three random variables: a uniform time-step $t\sim\mathcal{U}([0,1])$, samples from the base distribution ${\bm{z}}_{0}\sim P_{\mathcal{B}}$, and samples from the ground-truth conditional distribution ${\bm{z}}_{1}\sim P^{{\bm{z}}|{\bm{x}}}$.
We refer to Wildberger et al. (2024) for mathematical results on the relationship of $d_{\textrm{CFM}}$ and the (forward) Kullback-Leibler divergence. The hyperparameter $\sigma_{\textrm{min}}$, which is the variance at time $t=1$ in the Gaussian conditional probability paths, appears to have negligible influence when set to a sufficiently small value (Lipman et al., 2022). [^4] In our experiments, we follow Wildberger et al. (2024) and set $\sigma_{\textrm{min}}\vcentcolon=10^{-4}$ for all experiments.
@@
^$\hat{\mathcal{R}}_{\theta}=\sum_{i=1}^{N}\left\lvert\left\rvert v_{t^{(i)},{\bm{x}}^{(i)}}^{\theta}((1-(1-\sigma_{\textrm{min}})t^{(i)}){\bm{z}}_{0}^{(i)}+t^{(i)}{\bm{z}}_{1}^{(i)})-\frac{{\bm{z}}_{1}^{(i)}-(1-\sigma_{\textrm{min}}){\bm{z}}_{0}^{(i)}}{1-(1-\sigma_{\textrm{min}})t^{(i)}}\right\lvert\right\rvert_{2}^{2}.$ (6)
@@ 3.2 Sampling from the joint distribution
In order to learn a model that can perform posterior inference according to Section 3.1, we require to sample $({\bm{x}},{\bm{z}})\sim P^{{\bm{x}},{\bm{z}}}$. Given $p({\bm{x}},{\bm{z}})=p({\bm{x}}|{\bm{z}})p({\bm{z}})$, this is always possible as long as one can draw samples from $P^{{\bm{z}}}$ and then from $P^{{\bm{x}}|{\bm{z}}}$. Hence, this is a relatively weak requirement allowing for a broad variety of priors and observation models. More specifically, for ICL, we generate a training dataset $\mathcal{D}$ which comprises i.i.d. samples $\left\{\left({\bm{x}}^{(i)},{\bm{z}}^{(i)}\right)\right\}_{i=1}^{N}$ resulting from sampling ${\bm{z}}^{(i)}\sim P^{{\bm{z}}}$ and then ${\bm{x}}^{(i)}\sim P^{{\bm{x}}|{\bm{z}}^{(i)}}$. We use this simple yet fundamental and very general template to generate samples from the joint $P^{{\bm{x}},{\bm{z}}}$ for GLMs, factor analysis (FA), and Gaussian mixture models (GMMs) in our later applications.
GLM example
For example, assume that ${\bm{x}}\vcentcolon=({\bm{u}},y)$ is partitioned into covariates ${\bm{u}}$ and a response $y$ related via a conditional distribution $P^{y|{\bm{x}}}$ depending on a linear predictor ${\bm{u}}^{\top}\bm{\beta}$. This allows to define various GLM structures. In case of a fully Bayesian GLM, one further assumes a prior $P^{\bm{\beta}}$ on the regression coefficients and additionally on the variance $\sigma^{2}$ of the responses, which takes the role of a separate dispersion parameter, as well as a link function $g:\mathbb{R}\rightarrow\mathbb{R}$. Algorithm 1 specifies how a dataset $\mathcal{D}$ with i.i.d samples from $P^{{\bm{x}},{\bm{z}}}$ can be sampled in this case.
1 Initialize $\mathcal{D}\leftarrow\o$;
2 for $i=1,\ldots N$ do
3       draw $\bm{\beta}_{i}\sim P^{\bm{\beta}}$ ;
4       draw $\sigma^{2}_{i}\sim P^{\sigma^{2}}$ ;
5       for $j=1,\ldots,K$ do
6             draw $\bm{u}_{i,j}\sim P^{\bm{u}}$;
7             draw $y_{i,j}\sim p(y|g^{-1}\left(\bm{u}_{i,j}^{\top}\bm{\beta}_{i}\right),\sigma^{2}_{i})$;
8
9       end for
10      set ${\bm{x}}^{(i)}\vcentcolon=\left((\bm{u}_{i,j},y_{i,j})\right)_{j=1}^{K}$;
11       set ${\bm{z}}^{(i)}\vcentcolon=\bm{\beta}_{i}$;
12       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
13
14 end for
Algorithm 1 Generation of synthetic data for GLMs
Variations in the structure of the distributions $P^{\bm{\beta}}$, $P^{\sigma^{2}}$, $P^{y|{\bm{x}}}$, as well as $g$ give rise to different models. Examples include Bayesian ridge, Bayesian lasso and logistic regression (Box & Tiao, 2011; Murphy, 2023), which we all consider in our later experiments (see Section A.1 for details on the distributional setups and GLMs in general). Analogously, albeit with different data generating mechanisms, one can obtain samples for FA and GMMs, which we detail in Section A.2 and Section A.3.
3.3 Generating realistic data
GLM example continued.
In the aforementioned GLM case, the distribution of the covariates $P^{{\bm{u}}}$ does not affect the structure of $P^{{\bm{z}}|{\bm{x}}}$ in the data generating process (cf. Algorithm 1). We can therefore use a flexible prior $P^{{\bm{u}}}$ such as the TabPFN-“prior” (Hollmann et al., 2022) to generate covariates ${\bm{u}}$ and thereby effectively tackle the issue of model specification. More specifically, by generating a plethora of highly realistic samples of tabular covariates with different ranges, domains, and correlations, the in-context learner will learn the GLM structure on a broad mixture of covariate distributions.
@@ 3.4 The architecture
In order to implement the idea of learning full Bayesian inference in context, we extend ideas of diffusion transformers (Peebles & Xie, 2023), where the conditioning on the time $t$ is implemented via adaptive layer norm (adaLN) blocks initialized as the identity function. As we potentially require complex conditioning on the data ${\bm{x}}$, an additional transformer encoder is added. The input to the decoder is a vector in the form $1-(1-\sigma_{\textrm{min}})t){\bm{z}}_{0}+t{\bm{z}}_{1}$, which is treated as a sequence with length one and processed by a transformer decoder without self-attention, but the adaLN blocks. Therefore, the decoder has an equivalent interpretation as a multi-layer perceptron with skip-connections, cross-attention, and adaptive layer normalization. For the final processing in the decoder, only conditional feedforward layers with adaptive layer normalization are used. This corresponds exactly to the architecture of the decoder before, albeit without cross attention. We call this part an “MLP with Conditioning”. Samples for the time $t\in\left[0,1\right]$ are mapped onto a conditioning vector using several fully connected layers, which yields a richer representation of $t$ that is well-suited as an input to the adaLN blocks. Fig. 2 depicts of the resulting architecture.
@@
Figure 2: Architecture to perform ICL for full Bayesian inference.
@@ 3.5 Implementing Flow Matching
During the training phase, a tuple $({\bm{z}}_{1},{\bm{x}})$ is drawn from the distribution $P^{{\bm{z}},{\bm{x}}}$. Additionally, a time step $t\sim\mathcal{U}[0,1]$ and a sample ${\bm{z}}_{0}$ is drawn from the base distribution $P_{\mathcal{B}}$, which is a standard Gaussian for all our applications. Subsequently, the ground-truth conditional flow $\psi({\bm{z}}_{0}|{\bm{x}})=1-(1-\sigma_{\textrm{min}})t){\bm{z}}_{0}+t{\bm{z}}_{1}$ is computed, pushing forward $P_{\mathcal{B}}$ into $P^{{\bm{z}}|{\bm{x}}}$ up to time-point $t$. The transformer encoder processes ${\bm{x}}$ and the decoder takes the representation of the encoder into account in order to output $v_{t,{\bm{x}}}^{\theta}(\psi({\bm{z}}_{0}|{\bm{x}}))$. This output should match the vector field that describes how the ground-truth flow $\psi({\bm{z}}_{0}|{\bm{x}})$ continues at time $t$. The discrepancy to the ground-truth vector field is measured with the MSE-loss in Eq. 6.
In the sampling phase, we are given ${\bm{x}}$ and the goal is to sample from $P^{{\bm{z}}|{\bm{x}}}$. To do so, first a vector ${\bm{z}}_{0}\sim P_{\mathcal{B}}$ is drawn. The data ${\bm{x}}$ is passed through the encoder. The decoder defines a function that maps a time-point $t$ and a vector $\bm{\nu}$ onto a vector field: $(t,\bm{\nu})\mapsto v_{t,{\bm{x}}}^{\theta}(\bm{\nu})$ taking ${\bm{x}}$ into account. This function is given to an ODE-solver in order to forward-solve the corresponding ODE with boundary conditions $0\leq t\leq 1$.
@@ 4 Experiments
We use seven different scenarios for the GLMs, where we vary the prior distribution on the parameters, the conditional distribution of the response, and whether an intercept is included. For FA, we vary the form of the priors and dimensionalities of variables. For the GMMs, we investigate different dimensionalities as well as prior configurations. We refer to Appendix A for details on the model structure and scenarios.
@@
We evaluate the methods on 50 synthetic datasets and 17 real-world datasets from a benchmark suite proposed by Grinsztajn et al. (2022). We refer to Appendix J for more details on the preprocessing of the datasets.

### v2

@@ 3.1 Defining the Form of the Posterior
To learn the posterior distribution $P^{\bm{z}|\bm{x}}$ in context, we use the framework of flow matching (Lipman et al., 2022). More specifically, we utilize continuous normalizing flows (CNFs) to specify and ultimately sample from $P^{\bm{z}|\bm{x}}$. CNFs, currently excelling in the field of image synthesis (Esser et al., 2024), do not only allow to flexibly learn almost arbitrary distributions, but are also found to be more sample-efficient in training than for instance diffusion objectives (Lipman et al., 2022; Wildberger et al., 2024). Furthermore, unlike discrete normalizing flows (Papamakarios et al., 2021a), CNF objectives do not limit the architecture of the used neural network, allowing to incorporate complex conditioning on the data $\bm{x}$ in addition to flexibly modeling the posterior, which is a crucial aspect of our ICL framework. Refer to Appendix D for more information on CNFs.
@@ 3.1.1 Normalizing flows
Therefore, one assumes that samples from $P^{\bm{z}|\bm{x}}$ are generated by first drawing $\bm{z}^{(0)}\sim P_{\mathcal{B}}$, and then applying $\psi_{\theta}(\cdot|\bm{x})$, such that $\psi_{\theta}(\bm{z}^{(0)}|\bm{x})\sim P^{\bm{z}|\bm{x}}$. The base distribution $P_{\mathcal{B}}$ is commonly set to be a standard normal distribution, i.e., $P_{\mathcal{B}}=\mathcal{N}(0,I)$. The conditional flow $\psi_{\theta}(\cdot|\bm{x})$ is the object to be learned, such that our model of $P^{\bm{z}|\bm{x}}$ is defined as $Q^{\bm{z}|\bm{x}}_{\theta}\vcentcolon=[\psi_{\theta}(\cdot|\bm{x})]_{\sharp}P_{\mathcal{B}}$.
@@ 3.1.2 Continuous Normalizing Flows
$\frac{d}{dt}\psi_{\theta,t}(\bm{z}|\bm{x})=v^{\theta}_{t,\bm{x}}(\psi_{\theta,t}(\bm{z}|\bm{x})),\ \psi_{\theta,0}(\bm{z}|\bm{x})=\bm{z},$ (4)
where $0\leq t\leq 1$. The first condition $\frac{d}{dt}\psi_{\theta,t}(\bm{z}|\bm{x})=v^{\theta}_{t,\bm{x}}(\psi_{\theta,t}(\bm{z}|\bm{x}))$ means that $v^{\theta}_{t,\bm{x}}$ describes the change in $\psi_{\theta,t}(\bm{z}|\bm{x})$ at time $t$, and the second condition $\psi_{\theta,0}(\bm{z}|\bm{x})=\bm{z}$ implies that initially the flow is just the identity. The family of vector fields $v^{\theta}_{t,\bm{x}}$ is parameterized by a neural network whose parameters $\theta$ will be learned. In order to ultimately compute the flow $v^{\theta}_{1,\bm{x}}$, that yields $Q^{\bm{z}|\bm{x}}_{\theta}=[\psi_{\theta,1}(\cdot|\bm{x})]_{\sharp}P_{\mathcal{B}}$, a numerical ODE solver can be used to forward-solve the ODE, which ultimately corresponds to evaluating $\psi_{1,\bm{x}}$ at a data point $\bm{z}^{(0)}\sim P_{\mathcal{B}}$. This construction implies very generic assumptions regarding the structure of $Q^{\bm{z}|\bm{x}}$, which include the existence of a density of the target distribution wrt. the Lebesgue measure, and the assumption that $P^{\bm{z}|\bm{x}}$ can be represented by a mixture distribution over the marginal probability paths at time point $t=1$ (Lipman et al., 2022). Please note that the mathematical understanding of Flow Matching, its properties and assumptions, are still actively researched (Wildberger et al., 2024).
@@
$d_{\textrm{CFM}}\left(Q^{\bm{z}|\bm{x}}_{\theta},P^{\bm{z}|\bm{x}}\right)\vcentcolon=\\ \mathbb{E}\left[\left\lvert\left\rvert v_{t,\bm{x}}^{\theta}(\gamma_{t}(\bm{z}^{(1)}|\bm{z}^{(0)}))-{(\bm{z}^{(1)}-\omega\bm{z}^{(0)})}\right\lvert\right\rvert_{2}^{2}\right],$ (5)
where the expectation is taken w.r.t. to three random variables: a uniform time-step $t\sim\mathcal{U}([0,1])$, samples from the base distribution $\bm{z}^{(0)}\sim P_{\mathcal{B}}$, and samples from the ground-truth conditional distribution $\bm{z}^{(1)}\sim P^{\bm{z}|\bm{x}}$. We define $\gamma_{t}(\bm{z}^{(1)}|\bm{z}^{(0)})\vcentcolon=(1-{\omega}t)\bm{z}^{(0)}+t\bm{z}^{(1)}$.
We refer to (Wildberger et al., 2024) for mathematical results on the relationship of $d_{\textrm{CFM}}$ and the (forward) Kullback-Leibler divergence. The hyperparameter $\omega=1-\sigma_{\textrm{min}}$, where $\sigma_{\textrm{min}}$ is the variance at time $t=1$ in the Gaussian conditional probability paths, appears to have negligible influence when set to a value sufficiently close to one (Lipman et al., 2022). [^3] In our experiments, we follow (Wildberger et al., 2024) and set $\omega\vcentcolon=1-10^{-4}$ for all experiments.
In order to make optimizing
$\mathbb{E}_{\bm{x}\sim p(\bm{x})}\left[d_{\textrm{CFM}}\left(Q^{\bm{z}|\bm{x}}_{\theta},P^{\bm{z}|\bm{x}}\right)\right]$ (6)
@@
$\hat{\mathcal{R}}_{\theta}=\sum_{i=1}^{N}\left\lvert\left\rvert v_{t_{i},\bm{x}_{i}}^{\theta}(\gamma_{t_{i}}(\bm{z}_{i}^{(1)}|\bm{z}_{i}^{(0)}))+{\bm{z}_{i}^{(1)}-\omega\bm{z}_{i}^{(0)}}\right\lvert\right\rvert_{2}^{2}$ (7)
@@ 3.2 Sampling from the Joint Distribution
In order to learn a model that can perform posterior inference according to Section 3.1, we require to sample $(\bm{x},\bm{z})\sim P^{\bm{x},\bm{z}}$. Given $p(\bm{x},\bm{z})=p(\bm{x}|\bm{z})p(\bm{z})$, this is always possible as long as one can draw samples from $P^{\bm{z}}$ and then from $P^{\bm{x}|\bm{z}}$. Hence, this is a relatively weak requirement allowing for a broad variety of priors and observation models. More specifically, for ICL, we generate a training dataset $\mathcal{D}$ which comprises i.i.d. samples $\left\{\left(\bm{x}_{i},\bm{z}_{i}\right)\right\}_{i=1}^{N}$ resulting from sampling $\bm{z}_{i}\sim P^{\bm{z}}$ and then $\bm{x}_{i}\sim P^{\bm{x}|\bm{z}_{i}}$. We use this simple yet fundamental and very general template to generate samples from the joint $P^{\bm{x},\bm{z}}$ for GLMs, factor analysis (FA), and Gaussian mixture models (GMMs) in our later applications. Please refer to Appendix A for more details on the data generating processes.
@@ 3.3 The Architecture
In order to implement the idea of learning full Bayesian inference in context, we extend ideas of diffusion transformers (Peebles & Xie, 2023), where the conditioning on the time $t$ is implemented via adaptive layer norm (adaLN) blocks initialized as the identity function. As we potentially require complex conditioning on the data $\bm{x}$, an additional transformer encoder is added. The input to the decoder is a vector in the form $(1-{\omega}t)\bm{z}^{(0)}+t\bm{z}^{(1)}$, which is treated as a sequence with length one and processed by a transformer decoder without self-attention, but the adaLN blocks. Therefore, the decoder has an equivalent interpretation as a multi-layer perceptron with skip-connections, cross-attention, and adaptive layer normalization. For the final processing in the decoder, only conditional feedforward layers with adaptive layer normalization are used. This corresponds exactly to the architecture of the decoder before, albeit without cross attention. We call this part an “MLP with Conditioning”. Samples for the time $t\in\left[0,1\right]$ are mapped onto a conditioning vector using several fully connected layers, which yields a richer representation of $t$ that is well-suited as an input to the adaLN blocks. Figure 2 depicts of the resulting architecture.
@@
Figure 2: Architecture to perform ICL for full Bayesian inference. A relatively large transformer encoder, similar to that in TabPFN (Hollmann et al., 2022) processes a dataset $\bm{x}$ and yields a representation used in the decoder. The decoder outputs a vector field defining a flow for a given input vector conditioned on the encoder output and the time. We condition on the time in each cross-attention and each feed-forward block. Please note that the size of the parts of the architecture does not correspond to the number of allocated parameters in this figure.
@@ 3.4 Implementing Flow Matching
During the training phase, a tuple $(\bm{z}^{(1)},\bm{x})$ is drawn from the distribution $P^{\bm{z},\bm{x}}$. Additionally, a time step $t\sim\mathcal{U}[0,1]$ and a sample $\bm{z}^{(0)}$ is drawn from the base distribution $P_{\mathcal{B}}$, which is a standard Gaussian for all our applications. Subsequently, the ground-truth conditional flow $\psi(\bm{z}^{(0)}|\bm{x})=(1-{\omega}t)\bm{z}^{(0)}+t\bm{z}^{(1)}$ is computed, pushing forward $P_{\mathcal{B}}$ into $P^{\bm{z}|\bm{x}}$ up to time-point $t$. The transformer encoder processes $\bm{x}$ and the decoder takes the representation of the encoder into account in order to output $v_{t,\bm{x}}^{\theta}(\psi(\bm{z}^{(0)}|\bm{x}))$. This output should match the vector field that describes how the ground-truth flow $\psi(\bm{z}^{(0)}|\bm{x})$ continues at time $t$. The discrepancy to the ground-truth vector field is measured with the MSE-loss in Equation 7.
In the sampling phase, we are given $\bm{x}$ and the goal is to sample from $P^{\bm{z}|\bm{x}}$. To do so, first a vector $\bm{z}^{(0)}\sim P_{\mathcal{B}}$ is drawn. The data $\bm{x}$ is passed through the encoder. The decoder defines a function that maps a time-point $t$ and a vector $\bm{\nu}$ onto a vector field: $(t,\bm{\nu})\mapsto v_{t,\bm{x}}^{\theta}(\bm{\nu})$ taking $\bm{x}$ into account. This function is given to an ODE-solver in order to forward-solve the corresponding ODE with boundary conditions $0\leq t\leq 1$.
Table 1: Summarized results for GLMs. Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 7 different GLM scenarios. Comparison to the analytical solution when available and HMC otherwise. Lower is better for all metrics. The best average result is marked in bold.
Model Synthetic Real-World
C2ST MMD $\hskip 2.84544pt\mathcal{W}_{2}$ C2ST MMD $\mathcal{W}_{2}$
LA 1.000 2.770 2.049 1.000 2.091 0.849
VI: Diagonal 0.869 1.586 1.742 0.819 0.583 0.529
VI: Full 0.714 1.016 1.601 0.668 0.116 0.374
VI: Structured 0.711 0.929 1.580 0.664 0.109 0.370
VI: IAF 0.784 1.648 2.349 0.732 0.516 0.680
ICL (ours) 0.657 0.183 0.556 0.648 0.090 0.387
@@ 4 Experiments
We use seven different scenarios for the GLMs, where we vary the prior distribution on the parameters, the conditional distribution of the response, and whether an intercept is included. For FA, we vary the form of the priors and dimensionalities of variables leading to four different scenarios. For the GMMs, we investigate different dimensionalities as well as prior configurations also in four different scenarios. We refer to Appendix A for details on the model structure and scenarios.
For our experiments, we train a separate model from scratch for each GLM, GMM, and FA scenario using synthetic samples from the joint distribution $P^{\bm{x},\bm{x}}$; i.e. we train seven separate models to cover the GLM scenarios, six separate models for the FA scenarios and four separate models for the GMM scenarios. Please refer to Appendix B for more details.
@@
We evaluate the methods on 50 synthetic datasets and 17 real-world datasets from a benchmark suite for tabular regression problems proposed by Grinsztajn et al. (2022). We refer to Appendix C for more details on the preprocessing of the datasets.